If 400 ohm of resistance is made by adding four 100 ohm resistances of tolerance 5%, then the tolerance of the combination is

If 400 ohm
of resistance is made by adding four 100 ohm resistances of tolerance 5%, then the tolerance of the
combination is

Grade:11

3 Answers

Saurabh Koranglekar
3 years ago
Dear student

The answer is 5%

Regards
Arun
3 years ago
Each 100 ohm resistance have tolerance of 5%. i.e. 5 ohm.
Total tolerance of 4 such resistances is 4 × 5 = 20 ohms.
Total resistance of combination is 4 × 100 = 400 ohms.
Tolerance of the combination will be -
Tolerance = 20 / 400 × 100 %
Tolerance = 5 %
Therefore, tolerance of the combination will be 5 %.
Vikas TU
3 years ago
Total tolerance of 4 such resistances is 4 × 5 = 20 ohms. Total resistance of combination is 4 × 100 = 400 ohms. Therefore, tolerance of the combination will be 5 %.